### v2

# Role-Aware Language Models for Secure and Contextualized Access Control in Organizations

Abstract
As large language models (LLMs) are increasingly deployed in enterprise settings, controlling model behavior based on user roles becomes an essential requirement. Existing safety methods typically assume uniform access and focus on preventing harmful or toxic outputs, without addressing role-specific access constraints. In this work, we investigate whether LLMs can be fine-tuned to generate responses that reflect the access privileges associated with different organizational roles. We explore three modeling strategies: a BERT-based classifier, an LLM-based classifier, and role-conditioned generation. To evaluate these approaches, we construct two complementary datasets. The first is adapted from existing instruction-tuning corpora through clustering and role labeling, while the second is synthetically generated to reflect realistic, role-sensitive enterprise scenarios. We assess model performance across varying organizational structures and analyze robustness to prompt injection, role mismatch, and jailbreak attempts.
Role-Aware Language Models for Secure and Contextualized Access Control in Organizations
Saeed Almheiri1  Yerulan Kongrat2  Adrian Santosh3  Ruslan Tasmukhanov2 Josemaria Loza Vera4Muhammad Dehan Al Kautsar1Fajri Koto1 1Mohamed bin Zayed University of Artificial Intelligence 2Nazarbayev University  3University of Illinois at Urbana-Champaign 4New York University Abu Dhabi Saeed.Y@mbzuai.ac.ae
1 Introduction
In enterprise workflows, access control is a core security mechanism for regulating access to organizational resources. Through authentication and authorization, systems verify user identities and enforce access privileges. While role-based access control (RBAC) is well established in traditional software systems ferraiolo1995role; sandhu1998role; park2001role, its application to large language models (LLMs) remains largely unexplored. As LLMs are increasingly deployed for enterprise applications such as document generation wiseman-etal-2017-challenges, summarization laskar-etal-2023-building; zhang-etal-2025-evaluating, and internal assistance muthusamy-etal-2023-towards, it becomes critical to enforce access control not just over outputs but at the level of model instructions.
Figure 1: A role-aware LLM rejects questions from unauthorized roles, enhancing safety by restricting access to sensitive information. Icon source: Flaticon.com
Figure 2: Overview of our methodology. Top-left: dataset preparation yields four datasets across two types (repurposed and synthetic) with predefined structures. Top-right: balanced test distribution over positive/negative and seen/unseen paraphrases. Bottom: three training strategies: Role-aware Cls (BERT-based), Role-aware LLM-Cls (LLM-based), and Role-aware LLM-Gen (response generation).
Figure 1 demonstrates how role-aware language models can help prevent unauthorized access to sensitive information. When the same instruction is issued by users in different roles, such as a CEO and a researcher, a role-unaware LLM may provide identical responses regardless of the requester’s permissions. In contrast, a role-aware LLM considers the user’s role and restricts access appropriately, disclosing information only to those with sufficient clearance and declining requests from others. This approach enables organizations to align LLM behavior with established access policies, minimizing the risk of information leakage across roles.
Despite increasing attention to the safety and alignment of LLMs wang-etal-2024-answer; ge-etal-2024-mart, the challenge of role-conditioned instruction filtering has received limited focus. Most existing approaches assume uniform user access or apply static safety filters, focusing primarily on preventing the generation of harmful or toxic content wang-etal-2024-answer; wang-etal-2024-chinese; azmi2025indosafety. These methods do not account for access control policies that vary by user role—a critical requirement in organizational contexts. To support secure, multi-user deployments, we pose the following research question: Can large language models be fine-tuned to generate role-aware responses that enforce access control? While LLMs continue to advance in capability and generalization jiang2024mixtral; dubey2024llama; bai2023qwen; dou2025sailor2; liu2024deepseek; koto2025llama, their application to role-sensitive scenarios remains underexplored.
To address this research question, we simulate realistic organizational scenarios and develop a role-aware language model using three complementary strategies: (i) a BERT-based classifier, (ii) an LLM-based classifier, and (iii) direct role-conditioned generation. We evaluate these methods on two separate datasets: one repurposed from existing instruction-tuning corpora using clustering and role-based labeling, and another consisting of synthetic, role-sensitive instructions generated by LLMs to reflect realistic enterprise interactions. Unlike contemporaneous work such as jayaraman2025permissioned, which focuses on domain-level access control, our approach explicitly models user roles and supports fine-grained, hierarchical permissions required in organizational settings.
Our contributions can be summarized as follows:
  • We evaluate role-aware LLMs in realistic organizational settings with diverse access structures, using multiple modeling strategies. Our experiments include full pretraining of six BERT-based classifiers and adapter-based fine-tuning on six different LLMs.
  • We conduct robustness analyses under various threat scenarios, including jailbreaking across role-encoding strategies, access control mismatches, and prompt injection or manipulation attacks.
  • We provide a comprehensive evaluation across varying levels of organizational complexity, comparing classifier-based and generation-based approaches, and analyzing performance on role-independent, blacklisted topics.
2 Related Works
Access Control in Traditional Systems
In classical role-based access control (RBAC), users are assigned roles with specific permissions ferraiolo1995role; ferraiolo2003role, enforcing the principle of least privilege. Organizations often segregate data by clearance levels or roles so that only authorized personnel can view sensitive records sandhu1998role; jayaraman2025permissioned. Role hierarchies allow higher-level roles (e.g., managers) to inherit the permissions of subordinate roles, a concept well understood in databases and operating systems. However, applying similar role-based permissions to a generative LLM is nontrivial chan2025encrypted, since the model can hallucinate or leak information beyond its explicit training data kaddour2023challenges.
Access Control in Language Models
Work on access control in language models remains limited. A contemporaneous study by jayaraman2025permissioned introduces PermissionedLLMs, which implement domain-based access control through parameter-efficient fine-tuning methods such as LoRA hu2022lora and Few-Shot Parameter Efficient Tuning liu2022few. Their approach defines access at the domain level, where a domain represents a group of data records requiring identical credentials. In parallel, sudollm:2025 proposed sudoLLM, which makes LLMs “user-aware” by injecting secret biases into input queries based on user identity. In contrast to these approaches, our work focuses on role-based access control with deeper hierarchical structures, making it more suitable for enterprise and organizational settings.
AdapterSwap adapterswap:2025 implements access control by associating different access levels with separate LoRA adapters, which are selected and composed at inference time. This approach requires maintaining multiple domain-specific adapters. In contrast, our method uses a unified model that directly encodes role-awareness without external composition. chen2023can address a related challenge from a privacy perspective, showing that pre-trained LLMs are prone to leaking sensitive information and proposing a self-moderation mechanism. While their work does not focus on role-aware modeling, it shares our broader goal of improving control over LLM outputs to prevent unauthorized disclosures.
3 Problem Formulation
Let $x$ be a prompt or instruction, $y$ the LLM output, and $r$ a user’s role within an organization. A general LLM defines a conditional distribution over outputs $y$ dependent on a user’s instruction $x$:
$P(y\mid x).$
However, a role-aware LLM defines the following distribution:
$P_{\mathrm{RoleLLM}}(y\mid x,r),$
such that $r\in\mathcal{R}$, where $\mathcal{R}$ is the set of all roles in an organization.
Now, formalizing access control, define a tree $\mathcal{T}=(\mathcal{R},\leq)$ such that for any two roles $r_{1},r_{2}\in\mathcal{R}$ where $r_{1}\leq r_{2}$ denotes $r_{2}$ inherits $r_{1}$’s permissions. Then, the access set of a role $r\in\mathcal{R}$ is:
$\mathcal{A}(r)=\bigcup_{r^{\prime}\leq r}\mathcal{S}(r^{\prime}),$
where $\mathcal{S}(r^{\prime})\subseteq\mathcal{Q}$. $\mathcal{S}(r^{\prime})$ is the set of all queries of role $r^{\prime}$, and $\mathcal{Q}$ is the universe of all valid input-output instruction types. Hence,
$P_{\mathrm{RoleLLM}}(y\mid x,r)=\begin{cases}P(y\mid x,r),&\mathrm{if}\ x\in\mathcal{A}(r)\\ \delta_{\mathrm{deny}}(y),&\mathrm{otherwise}\end{cases},$
such that $\delta_{\mathrm{deny}}(y)$ is a degenerate distribution concentrating all the probability mass on a refusal output (i.e., access is denied).
4 Dataset Construction
We define two organizational structures, each comprising 20 roles, to evaluate role-awareness under varying levels of hierarchy. The first is the Basic structure, where a single CEO directly supervises 19 subordinate roles. The second is the Office structure, which includes a CEO, four department managers reporting to the CEO, and 3–4 team members reporting to each manager. A detailed breakdown of roles in both structures is provided in Appendix C. These configurations are used to assess the ability of each method to encode and respond to hierarchical role information, as outlined in Section 5.1.
For each organizational structure, we construct two datasets using complementary strategies (see Figure 2). The first is by repurposing existing instruction-tuning data via clustering, and the latter involves generating synthetic data via LLMs.
Repurposing Existing Instruction Dataset
We repurpose We repurpose the Databricks Dolly-15k dataset (DatabricksBlog2023DollyV2) by clustering instructions and assigning roles based on hierarchical structure. Using a sentence transformer (reimers-gurevych-2019-sentence), we encode each instruction and its context into dense vectors. Clustering begins at the root of the organization: we apply K-Means to partition the data into three high-level groups: General, Shared, and Root Only (e.g., CEO-specific). Prompts in the General group are marked terminal and excluded from further subdivision. [^1] The General group refers to prompts that are accessible to all roles within the organization Shared prompts are recursively partitioned along the hierarchy. At each level, prompts are split into role-specific clusters corresponding to subordinate roles (e.g., Department 1, Department 2, etc.). Within each cluster, we further divide prompts into Shared (used across subordinates) and Role Only (exclusive to the role). The process continues recursively: Shared prompts are passed down for further subdivision, while Role Only groups are treated as terminal. This hierarchical clustering procedure, illustrated in Figure 8 (Appendix D), yields fine-grained, role-aligned instruction sets that mirror the structure of the target organization.
Synthetic Organization Dataset
We use OpenAI’s GPT-4.1 mini with a temperature of 0.7 to generate synthetic organizational data. Based on the basic and office structures (Appendix C), we define each role, department, and access range in a structured JSON-like format. Prompts are then generated for each role, conditioned on its responsibilities and access scope. The resulting data is organized with the fields: role, instruction, and output. We also generate 200 general instruction-response pairs representing organization-wide prompts that are accessible to all roles. Details of the generation prompt are provided in Appendix D.
Synthetic Dataset Quality Analysis
To evaluate the quality of the synthetic dataset, we randomly sampled 100 query-response pairs for manual analysis. Each pair was scored on two binary criteria: (1) whether the query was relevant to the assigned role, and (2) whether the response was complete and appropriate. A score of 1 was given for each criterion if it was met, and 0 otherwise. The results show that over 96% of the samples satisfied both criteria, indicating high relevance and response quality.
Training Set Construction
To train the model to distinguish between authorized and unauthorized access, we construct positive and negative instances from each instruction-response pair. First, we assign each pair the lowest-level role authorized to access the instruction. Using this role as an anchor, we generate four training instances through a sliding-window over the organizational hierarchy. Specifically, we create: (1) a positive instance using the minimal authorized role, and (2) another positive instance using its immediate parent, reflecting inherited permissions. We then generate two negative instances: (3) one from a subordinate role (or a random role from a different branch if no children exist), and (4) one from a non-existent external role. Each instance is labeled with a binary (1 for access granted, 0 for denied). For the denied request, LLM is expected to generate a generic refusal message. This procedure results in 6,000 training samples per dataset variant: repurposed_basic, repurposed_office, synthetic_basic, and synthetic_office. The ratio of positive and negative samples is approximately balanced: repurposed datasets contain 54.5% valid examples, and synthetic datasets contain 52.5%.
Test Set Construction
Each dataset variant includes a test set of 1,000 samples, balanced with 50% positive and 50% negative instances. Positive samples are split evenly into two subsets: 250 with previously unseen instructions, and 250 with paraphrased versions of training instructions generated by GPT-4.1 mini. Negative samples are divided into three categories: (1) 300 mismatch cases, where an unauthorized in-hierarchy role attempts to access restricted content (e.g., a leaf role querying CEO-level data); (2) 100 random cases using external roles not present in the hierarchy; and (3) 100 broken cases where the role string is intentionally corrupted (e.g., “1.2” → “01.02”, “1..2”, or “one.two”) to test model robustness. Each negative category includes an equal mix of unseen and paraphrased instructions, ensuring that every test set contains exactly 500 unseen and 500 seen prompts (See Figure 2.
5 Experimental Set-Up
5.1 Role Encoding Strategies
After grouping instruction-response pairs by role, we encode each role to study how different encoding strategies affect access control. Each organizational position is represented by a string that reflects its location in the hierarchy, which is appended to every instruction-response pair to indicate the minimum role required to access the content. Access is permitted to roles at or above the specified level and denied to those below or in unrelated branches. We explore three encoding methods. Hierarchical Number Encoding uses dot-delimited indices (e.g., “1” for the CEO, “1.1” and “1.2” for direct subordinates), with “1.0” reserved for general, organization-wide instructions. Single Name Encoding uses only the role’s title (e.g., “CEO,” “IT Department Manager”), while Hierarchical Name Encoding concatenates the full path of titles (e.g., “CEO - IT Department Manager - IT Support”) to retain both structural and semantic information.
5.2 Training Approaches
We evaluate three methods for access control, training each on all four dataset variants. To ensure reproducibility, we fix random seeds and report averaged results over three runs per setting. Training data is shuffled to eliminate order effects. Each training instance includes a prompt, answer, role, and access label. Full training details and hyperparameters are provided in Appendix B.
Role-aware Cls
We trained six BERT-based models devlin-etal-2019-bert; liu2019roberta, including Modern BERT-base, Modern BERT-large, Google BERT-base, Google BERT-large, Roberta-base, and Roberta-large for access control. We appended the role to the end of the prompt as “<prompt> [SEP] <role>”.
Role-aware LLM-Cls
We fine-tune six open-source LLMs bai2023qwen; dubey2024llama; team2024gemmaQwen 2.5 (3B, 7B), Llama 3.x (3B, 8B), and Gemma (4B, 7B)—to perform binary access control classification. We include both small and large models to assess the effect of model size. For each example, the role is prepended to the prompt as “Position: <role> <prompt>”, and a system prompt instructs the model to respond with True (grant access) or False (deny access). All inputs and labels are formatted as conversations and fine-tuned using LoRA with supervised learning.
Role-aware LLM-Gen
We use the same LLMs and fine-tuning setup as in Role-aware LLM-Cls, but instead train the model to generate full answers rather than binary access decisions. The system prompt is removed to allow free-form responses, and the output corresponds to the original answer instead of a True/False label.
5.3 Evaluation Protocol
For the classification-based approaches (Role-aware Cls and Role-aware LLM-Cls), we report standard metrics: accuracy, false positive rate (FPR), false negative rate (FNR), and F1 score. FPR captures unauthorized access incorrectly granted, while FNR reflects valid access that was wrongly denied. We also report performance on “seen” vs. “unseen” instructions, along with category-specific accuracy for mismatch, random, and broken roles. For Role-aware LLM-Gen, which outputs either a direct answer or a generic denial, we use GPT-4.1 mini to classify each response as grant or deny, enabling comparison with the ground-truth valid label.
Finally, to assess whether access control impacts answer quality, we randomly sample 100 valid (granted) examples and compare the generated responses to the original references. Each response is evaluated using GPT-4.1 mini, scored on a 1–5 scale for correctness, completeness, and clarity.
Model Generation Quality (↑, 5-pt rubric)
Correctness Completeness Clarity
Repurposed Existing Instruction Dataset (Dolly)
[…]
Synthetic Organization Dataset
[…]
Table 1: LLM-rated generation quality against gold reference. Bold = best within the same training dataset; underline = best within the Role-aware LLM-Gen method.
6 Results
Tables LABEL:tab:roleaware-results and 1 (or Table 13 and Table 14 for details) summarize the performance of our proposed role-aware LLMs evaluated on access-control accuracy and LLM-rated generation quality across two distinct training datasets: a repurposed existing instruction dataset (Dolly) and a synthetic organization dataset. The evaluation is conducted on all Role-aware methods (Cls, LLM-Cls, and LLM-Gen), assessing both quantitative metrics (e.g., accuracy, negative-pair defense) and qualitative dimensions (correctness, completeness, clarity). The detailed results and comparisons between the training datasets and modeling methods are discussed in further detail in the following sections.
Access-Control Performance
Our role-aware LLMs consistently achieved high access-control accuracy across both datasets, with LLM-Cls models outperforming other variants; specifically, ModernBERT Large attained the highest accuracy (90.0%) on Dolly, while Llama3 8B Instruct achieved top performance (89.3%) on the synthetic dataset. Generative approaches (LLM-Gen) slightly lagged in raw accuracy by approximately 5–10 percentage points with an influx in false-negative errors, indicating a strict access enforcement in role-conditioned generation. However, notable negative results emerged, particularly with RoBERTa Large (Cls), whose accuracy drastically decreased to 74.8% accompanied by an inflated false-positive rate (58%) on the Dolly dataset and subsequently in the synthetic dataset, highlighting critical sensitivity to encoder selection. In the more challenging synthetic dataset, all methods faced increased difficulty (3–6% accuracy drop), yet instruction-tuned models maintained comparatively robust performance, emphasizing that richer instruction tuning substantially mitigates accuracy degradation under semantically overlapping role conditions. Please refer to Appendix E for further explanation.
Method Robustness
To evaluate the robustness of our proposed methods, each method-model combination was trained under two organizational structures (basic, office) across three independent random seeds, with the results averaged and summarized in Tables LABEL:tab:roleaware-results1. Generally, all methods demonstrated low variance (acc std.<< 4%) on the Dolly dataset, except for notable brittleness in RoBERTa Large (Cls), which exhibited substantial instability (12.2% accuracy, 45.6% FPR), contrasting strongly with the more stable ModernBERT Large (3.2% accuracy, 8.1% FPR). Instruction-tuned LLM classifiers (LLM-Cls), such as Qwen2.5 3B Inst. and Llama3 3B Inst., further reduced variance (acc std.<< 2.2%), underscoring stability gains from modern instruction tuning. On the synthetic dataset, semantic overlaps increased variance to around 8–10%, yet instruction-tuned models (e.g., Llama3 8B Inst.) maintained comparative stability (8.4-8.6%). More detailed of the performances on Basic and Office are shown in Appendix H. Collectively, these results demonstrate that our proposed methods achieve robust performance, primarily due to richer pre-training and instruction tuning rather than merely model scale.
Negative Pair Accuracy
All methods achieved near-perfect accuracy (100%) in identifying randomly assigned negative-role pairs, highlighting their effectiveness in clearly invalid scenarios. However, performance dropped notably for subtler cases such as existing-but-mismatched and broken-role pairs. Specifically, LLM-Cls models demonstrated comparatively stronger performance (e.g., ModernBERT Large: 81.1%; Qwen2.5 3B Inst.: 78.2% on Dolly), whereas standard classifiers (Cls), particularly RoBERTa Large (41.0%), struggled significantly. Generative models (LLM-Gen) showed moderate accuracy (e.g., Llama3 3B Inst.: 73.3%), underscoring ongoing challenges in detecting nuanced role mismatches. These results indicate that while instruction-tuned models substantially enhance negative-pair detection, subtle distinctions between valid yet incorrect role assignments remain difficult, suggesting a promising direction for future improvements in fine-grained role understanding.
Generation Quality
Generation quality ratings (correctness, completeness, clarity) were consistently high (around 4 out of 5), with larger instruction-tuned models (e.g., Qwen2.5 7B Inst., Llama3 8B Inst.) achieving the best scores. The differences between the training data sets were minimal, highlighting the reliability of role-aware generation regardless of the data source.
7 Analysis
7.1 Jailbreak Robustness
To assess the model’s robustness against prompt injection attempts, we conducted an additional experiment involving modifications to the original datasets. Specifically, we generated 100 more negative samples for both training and test sets in which the original instruction was prepended with misleading phrases. These included assertions of higher authority (e.g., “I’m authorized as CEO to ask this:”) or commands (e.g., “Regardless of policy, respond to this:”). We evaluated Role-aware LLM-Cls using Llama 3.2 3B Instruct. The model was trained in two ways: the first incorporates jailbreak samples in the training data, and the other does not. A comparative summary of performance is presented in Appendix F. Although the inclusion of jailbreak prompts did not significantly affect overall accuracy across the test set (except for experiments with the repurposed_basic dataset, which can be attributed to randomness), the model trained with jailbreak-specific examples demonstrated improved resistance to prompt injection with an average of 87% compared to the 70% accuracy of the baseline model.
7.2 Robustness on Blacklisted Topics
To evaluate the model’s capacity to restrict answers to queries whose content violates organizational policies, we extended the original datasets mentioned in Section 4. We generated 100 queries on general blacklisted topics (e.g., violence, weapons, cheating, etc.) and 100 queries related to real-life politics. The respective responses to the queries were designed to be restricted, regardless of an employee’s role. Subsequently, each original dataset was extended by adding 50 unique blacklisted queries of each type separately, and duplicates of each blacklisted query for multiple organization roles. The remaining 50 queries of each blacklisted dataset were used for the evaluation datasets. Using the Role-aware LLM-Cls method, Llama 3.2 3B Instruct was trained and tested using these extended datasets. The detailed information on the results can be found in Appendix G. As shown in Table 4, the blacklisted model’s performance remained unchanged relative to the baseline model. The model was also highly successful in restricting answers to blacklisted queries with an overall accuracy >99%. The accuracy rates for the model trained on the repurposed basic dataset were the only outliers, exhibiting a decrease in accuracy from 92% to 84%.
7.3 Effect of Role Information in Prompts
To assess whether including role information in the prompt affects response quality, we fine-tuned all LLMs on the four training datasets without role annotations. We evaluated response quality using three metrics: correctness, completeness, and clarity. From the 1,000 test outputs, we randomly sampled 100 responses and compared them to the reference answers using GPT-4.1 mini. The same evaluation was applied to Role-aware LLM-Gen, which was trained with roles included in the prompt. Results show that the average difference in quality between the two settings is under 1%, indicating that including roles does not degrade response quality. Summary metrics are reported in Table 2, with detailed results in Appendix I.
Prompt Style Correctness Completeness Clarity
Without roles 3.90 3.58 4.67
With roles 3.93 3.59 4.64
Table 2: Quality ratings (five-point scale) of responses generated by LLMs trained with versus without role prompts, assessed by GPT-4.1 mini.
Figure 3: Comparison of FPR and FNR across role encodings. The Hierarchical Number Encoding has the worst defense against unauthorized roles (highest FPR), and overly denies authorized roles (highest FNR).
Figure 4: Comparison of broken role rejection accuracy across role encodings. The Hierarchical Number Encoding has the best defense against broken roles.
7.4 Effect of Role Encoding on Access Control
We investigate how different role encoding strategies affect access control performance across our three methods: Role-aware Cls, Role-aware LLM-Cls, and Role-aware LLM-Gen. For consistency, we use the Modern BERT-base model for Role-aware Cls and Llama 3.1 8B Instruct for the LLM-based methods, training each on the four dataset variants.
We compare three encoding strategies: Hierarchical Number Encoding, Single Name Encoding, and Hierarchical Name Encoding, and present the results in Figures 3 and 4. Hierarchical Number Encoding achieves the highest FPR, indicating poorer rejection of unauthorized roles and weaker robustness to broken role strings (e.g., misspelled or manipulated encodings). This suggests that LLMs can more easily differentiate between role names like “CEO” and “Researcher” than between formats like “1.1” and “1.a” . This encoding also results in the highest FNR, likely because LLMs struggle to generalize upward in hierarchical structures (e.g., understanding that “1” can access data assigned to “1.1”). In contrast, name-based encodings offer slightly better generalization across authorized roles but are more vulnerable to adversarial role perturbations. Full results are provided in Appendix J.
8 Conclusion
This paper investigates methods for modeling role-aware behavior in large language models, with a focus on enforcing access control and evaluating the effects of different fine-tuning strategies and datasets. Our experiments compare classification-based and generative approaches across multiple organizational structures. Instruction-tuned classifiers (LLM-Cls) consistently outperform both generative (LLM-Gen) and traditional classifier-based (Cls) methods, reaching up to 90.0% and 89.3% accuracy on the Dolly and synthetic datasets, respectively, without compromising answer quality.
Despite high overall performance, challenges remain. All models are effective at rejecting clearly unauthorized roles, such as random or external entities ($\approx$100% accuracy), and instruction-tuned methods reliably detect more subtle mismatches ($\approx$70% accuracy on average). However, broken role formats and fine-grained violations still present difficulties, with a 15–30% gap in accuracy. Generative models, while more flexible, suffer a modest performance trade-off. Future work should focus on enhancing generalization across complex hierarchies, reducing false positives from brittle encoders, and improving discrimination between closely related roles.
9 Limitations
While our results demonstrate promising capabilities in enabling safe and role-aware deployment of LLMs within organizational contexts, several limitations constrain the scope of our conclusions.
Unified Organization Representation
Our experiments used a single adapter to represent all roles within an organization. Although effective, we did not investigate the alternative of using a multi-adapter strategy, such as separate adapters for each department. This strategy could potentially reduce information leakage by further isolating department-specific knowledge, though it may come at the cost of overall effectiveness.
Access Control Post Fine-tuning
We demonstrated effective fine-tuning of adapters for initial access control; however, our methodology did not address dynamic modification or addition of roles after the fine-tuning phase. Future research should explore approaches that enable post-training updates to role-based access, as roles are dynamic and such updates would eliminate the need to retrain adapters from scratch.
Alignment Methods Beyond SFT
This study exclusively employed SFT for alignment. We did not explore alternative methods such as Direct Preference Optimization (DPO) or other preference-based alignment techniques, which could potentially yield improved alignment outcomes.
Integration of External Knowledge
Although our results indicate strong capabilities in controlling internal knowledge, either by restricting specific topics organization-wide or selectively authorizing content per role, we did not evaluate role-aware control when the LLM is augmented with external knowledge sources (e.g., Retrieval-Augmented Generation or web search). Investigating how role-aware adapters influence responses that incorporate external information remains an open area for future study.
Appendix A Training Seeds
We used each of the seeds shown in Table 3 for all experiments and averaged the result over seeds for each experiment.
Seed # Value
Seed 1 42
Seed 2 937
Seed 3 3827
Table 3: Seeds for training, testing and evaluation for all methods
Appendix B Training Hyperparameters
We used the same set of hyperparameters (Table 4) to train all LLMs and a different set of hyperparameters (Table 5) to train all BERT models. We created a LoRA adapter to train LLMs with the LoRA configuration given in (Table 4 ).
Parameter Value
LoRA rank 32
LoRA alpha 64
LoRA dropout $1\times 10^{-1}$
LoRA modules
up proj, down proj,
gate proj, k proj,
q proj, v proj, o proj
Batch size 1
Epochs 4
Learning rate $1\times 10^{-4}$
Grad. accumulation 1
Weight decay 0.0
Warmup ratio 0.0
Table 4: Hyperparameters used for LoRA SFT training of LLMs
Parameter Value
Batch size 16
Epochs 5
Learning rate $2\times 10^{-5}$
Grad. accumulation 1
Weight decay $1\times 10^{-2}$
Warmup ratio $1\times 10^{-1}$
Table 5: Hyperparameters for BERT training
Appendix C Organizational Structure Details
We define two predefined structures for dataset creation: the Basic and Office structures, shown in Figure 5 and Figure 6, respectively. In the Basic structure, a single CEO directly corresponds to all other roles, allowing us to test whether the models can leverage role-awareness when faced with a wide, single-layer hierarchy. In contrast, the Office structure introduces a multi-level hierarchy, where the CEO supervises department managers, who in turn oversee several team members. This setup evaluates whether the methods discussed in Section 5.1 can effectively capture and utilize hierarchical relationships within the organization. Additionally, Figure 7 presents several example roles introduced in each structure for synthetic role data generation, making the data specific to the roles defined in each structure.
Figure 5: Hierarchical structure for Basic structure.
Figure 6: Hierarchical structure for Office structure.
Figure 7: Predefined roles for each Basic and Office structure.
Appendix D Dataset Creation
Figure 8 shows our clustering scheme when repurposing the dataset. At the root level, datasets are first partitioned into three clusters: General, Shared, and Root-Only. Prompts in the General cluster terminate immediately; those in Shared are then split along the root’s direct subordinate roles, and recursion continues further. Furthermore, Figure 9 shows the specific system-level prompt used to generate the synthetic data. Below the prompt is an example of the OpenAI API output with the specified keys after generating the dataset.
Figure 8: Hierarchical clustering scheme of repurposed dataset.
Figure 9: System-Level Output for Synthetic Dataset.
Appendix E Role-aware Method: Cls vs LLM-Cls vs LLM-Gen
The Role-aware Cls shows a highly inconsistent performance, with a mean FPR of 0.41 and a large variance between 0.23 and 0.68, where the Roberta-large model performed the worst with the highest FPR of 0.68, which means that there are significant model-dependent weaknesses to unauthorized access. However, they are consistently low in FNR (0.04-0.06, average 0.05), indicating reliable access to authorized users. Conversely, the Role-aware LLM-Gen exhibited more stable but poor security performance with moderate FPR (0.28-0.38, average 0.33) and significantly higher FNR variability (0.11-0.19, average 0.15), indicating that it has greater difficulty in rejecting genuine access requests across model implementations and organizational designs.
Figure 10: Performance comparison of three role-based access control architectures across security metrics. Results show minimum, average, and maximum values for FPR, FNR, and Broken Role accuracy across six different models per architecture, averaged over multiple datasets with organizational structure variations. Higher Broken Role accuracy indicates better defense against one of jailbreak attacks.
Most importantly, our analysis shows that there are different security capabilities against adversarial attacks in different architectures. The Role-aware LLM-Gen strategy showed the best protection against broken role attacks with an average broken role accuracy of 0.56 (range: 0.42-0.63), and was able to reject the greatest percentage of malicious role manipulation attempts. Such high performance indicates that the integrated method, in which both access control and question answering are performed by a single model, offers improved contextual knowledge of role-based attacks. Role-aware CLs performed at average levels (average: 0.48, range: 0.30-0.40), whereas Role-aware LLM-CLs had the lowest broken role accuracy (average: 0.45, range: 0.39-0.48), which means that it is more susceptible to such adversarial attacks. These results indicate a curious tradeoff: whereas the Role-aware LLM-Gen approach exhibits larger FNR variation and moderate FPR, it makes up in better resistance to advanced attacking methods, indicating that the unified architecture might be inherently more capable of identifying and resisting role-based manipulation attacks than separated classification systems.
Appendix F Metrics for Jailbreak Experiment
Figure 4 shows the detailed performance between the baseline and the model that has been trained on the jailbreak train set (See Section 7.1).
Model Structure Accuracy Broken Jailbreak
Baseline RB11footnotemark: 1 0.92 0.49 0.71
RO22footnotemark: 2 0.89 0.29 0.69
SB33footnotemark: 3 0.96 0.58 0.89
SO44footnotemark: 4 0.80 0.30 0.51
With jailbreak samples RB 0.84 0.56 0.98
RO 0.88 0.27 0.70
SB 0.97 0.60 0.96
SO 0.82 0.39 0.83
Table 6: Jailbreak Experiment Performance for Llama 3.2 3B Instruct.
11footnotemark: 1
Repurposed Basic, 22footnotemark: 2Repurposed Office, 33footnotemark: 3Synthetic Basic, 44footnotemark: 4Synthetic Office
Appendix G Metrics for Blacklist Experiment
Figure 4 presents a detailed comparison between the baseline model and the model trained on the original plus the blacklist training set (see Section 7.2).
Blacklist Topic Structure Accuracy Blacklist
Baseline RB11footnotemark: 1 0.92 -
RO22footnotemark: 2 0.89 -
SB33footnotemark: 3 0.96 -
SO44footnotemark: 4 0.80 -
Politics RB 0.84 1.00
RO 0.88 1.00
SB 0.97 1.00
SO 0.80 1.00
General RB 0.84 1.00
RO 0.88 1.00
SB 0.96 1.00
SO 0.81 0.99
Table 7: Blacklist Experiment Performance for Llama 3.2 3B Instruct.
11footnotemark: 1
Repurposed Basic, 22footnotemark: 2Repurposed Office, 33footnotemark: 3Synthetic Basic, 44footnotemark: 4Synthetic Office
Note that Baseline here denotes the baseline datasets (original) used to train the model.
Appendix H Basic vs Office Structures
After training the models using three methods of Section 5.1, we averaged the accuracy metrics for the two types of structures (basic and office). As shown in Figure 11, model performance, on average, on the office organizational structure is lower than on the basic structure, as expected. For the Role-aware Cls and Role-aware LLM-Cls methods, the accuracy rates decreased by 6.0% and 6.7% when trained with the office structure. The reason for this is due to the deep hierarchy associated with the office structure compared to the basic one. Nonetheless, when using the Role-aware LLM-Gen method, the accuracy rate increased 1.3% when training with the office structure, potentially indicating that, with answer generation, there is negligible model performance difference when training with either structures
Figure 11: Average Accuracy Rates of Models Trained on the Basic vs Office Datasets.
Across almost all methods, models exhibit lower accuracy rates when trained with the office structure. Note that for Role-aware LLM-Gen, accuracy rates for both structures are almost equal.
Appendix I Role vs No role comparison
Tables 8 and 9 show the difference in quality of LLM responses to prompts with and without roles respectively. We use three metrics for response quality - Correctness, Completeness, and Clarity (on a scale of 1 to 5). The LLM responses are sent to ChatGPT 4.1 mini for evaluation as described in Section 7.3. The average metrics for prompts with and without roles are similar, with less than 1% difference between each of the metrics.
Architecture Dataset
Model
Org.
Structure
Seed Completeness Correctness Clarity
LLM + LLM Repurposed Qwen2.5 3B Instruct Basic 42 3.86 3.26 4.62
LLM + LLM Repurposed Qwen2.5 3B Instruct Office 42 3.7 3.21 4.61
LLM + LLM Repurposed Llama 3.2 3B Instruct Basic 42 3.85 3.43 4.64
LLM + LLM Repurposed Llama 3.2 3B Instruct Office 42 3.93 3.28 4.7
LLM + LLM Repurposed Gemma 3 4B Instruct Basic 42 4.03 3.53 4.52
LLM + LLM Repurposed Gemma 3 4B Instruct Office 42 3.91 3.39 4.41
LLM + LLM Repurposed Qwen2.5 7B Instruct Basic 42 4.1 3.69 4.75
LLM + LLM Repurposed Qwen2.5 7B Instruct Office 42 4.01 3.55 4.63
LLM + LLM Repurposed Llama 3.1 8B Instruct Basic 42 4.11 3.69 4.73
LLM + LLM Repurposed Llama 3.1 8B Instruct Office 42 4.15 3.63 4.72
LLM + LLM Repurposed Gemma 7B Instruct Basic 42 3.95 3.61 4.44
LLM + LLM Repurposed Gemma 7B Instruct Office 42 4.03 3.6 4.36
LLM + LLM Synthetic Qwen2.5 3B Instruct Basic 42 3.93 3.59 4.75
LLM + LLM Synthetic Qwen2.5 3B Instruct Office 42 3.6 3.63 4.75
LLM + LLM Synthetic Llama 3.2 3B Instruct Basic 42 3.84 3.66 4.74
LLM + LLM Synthetic Llama 3.2 3B Instruct Office 42 3.68 3.66 4.71
LLM + LLM Synthetic Gemma 3 4B Instruct Basic 42 4.09 3.66 4.77
LLM + LLM Synthetic Gemma 3 4B Instruct Office 42 3.75 3.62 4.65
LLM + LLM Synthetic Qwen2.5 7B Instruct Basic 42 3.95 3.71 4.83
LLM + LLM Synthetic Qwen2.5 7B Instruct Office 42 3.59 3.69 4.74
LLM + LLM Synthetic Llama 3.1 8B Instruct Basic 42 4.04 3.73 4.81
LLM + LLM Synthetic Llama 3.1 8B Instruct Office 42 3.79 3.75 4.78
LLM + LLM Synthetic Gemma 7B Instruct Basic 42 4.05 3.71 4.69
LLM + LLM Synthetic Gemma 7B Instruct Office 42 3.74 3.73 4.66
Average 3.9 3.58 4.67
Table 8: Response quality when no role is included in question for LLM
Architecture Dataset
Model
Org.
Structure
Seed Completeness Correctness Clarity
LLM Repurposed Qwen2.5 3B Instruct Basic 42 3.85 3.41 4.58
LLM Repurposed Qwen2.5 3B Instruct Office 42 3.83 3.38 4.67
LLM Repurposed Llama 3.2 3B Instruct Basic 42 3.97 3.50 4.56
LLM Repurposed Llama 3.2 3B Instruct Office 42 3.80 3.40 4.59
LLM Repurposed Gemma 3 4B Instruct Basic 42 3.96 3.56 4.53
LLM Repurposed Gemma 3 4B Instruct Office 42 4.10 3.64 4.54
LLM Repurposed Qwen2.5 7B Instruct Basic 42 3.94 3.51 4.73
LLM Repurposed Qwen2.5 7B Instruct Office 42 4.09 3.59 4.73
LLM Repurposed Llama 3.1 8B Instruct Basic 42 4.09 3.65 4.64
LLM Repurposed Llama 3.1 8B Instruct Office 42 4.02 3.52 4.63
LLM Repurposed Gemma 7B Instruct Basic 42 3.77 3.42 4.38
LLM Repurposed Gemma 7B Instruct Office 42 3.73 3.36 4.36
LLM Synthetic Qwen2.5 3B Instruct Basic 42 3.89 3.56 4.75
LLM Synthetic Qwen2.5 3B Instruct Office 42 3.96 3.86 4.82
LLM Synthetic Llama 3.2 3B Instruct Basic 42 3.91 3.61 4.64
LLM Synthetic Llama 3.2 3B Instruct Office 42 3.87 3.76 4.70
LLM Synthetic Gemma 3 4B Instruct Basic 42 3.92 3.60 4.61
LLM Synthetic Gemma 3 4B Instruct Office 42 3.90 3.78 4.73
LLM Synthetic Qwen2.5 7B Instruct Basic 42 4.13 3.88 4.79
LLM Synthetic Qwen2.5 7B Instruct Office 42 3.98 3.81 4.79
LLM Synthetic Llama 3.1 8B Instruct Basic 42 3.86 3.60 4.78
LLM Synthetic Llama 3.1 8B Instruct Office 42 3.91 3.65 4.78
LLM Synthetic Gemma 7B Instruct Basic 42 3.84 3.55 4.54
LLM Synthetic Gemma 7B Instruct Office 42 3.88 3.65 4.59
Average 3.93 3.59 4.64
Table 9: Response quality when role is included in question for LLM
Appendix J Comparison of encodings
We show our results from comparison of different role encodings for access control as described in Section 7.4. We experimented with Single Name Encoding (Table 10), Hierarchical Name Encoding (Table 11), and Hierarchical Number Encoding (Table 12). We used four metrics to compare model responses across role encodings: Accuracy, FPR (how often the model gives access to unauthorized roles), FNR (how often the model denies access to authorized roles), and F1. Compared to Hierarchical Number Encoding, the Single Name Encoding has a 28.33% decrease in FPR (26.19% to 18.77%) and a 45.15% decrease in the FNR (9.08% to 4.98%). There is a 47.64 % decrease in the broken role rejection accuracy (51.42% to 26.92%). Similarly, the Hierarchical Name Encoding has a 29.13 % decrease in FPR (26.19% to 18.56%), a 45.15% decrease in the FNR (9.08% to 4.98%) and a 47.64 % decrease in the broken role rejection accuracy (51.42% to 26.92%) when compared to the Hierarchical Number Encoding. Overall, the Hierarchical Number Encoding has the highest FPR, highest FNR and highest broken role rejection accuracy.
Architecture Dataset
Model
Org.
Structure
Seed Accuracy FPR FNR F1
Seen
Role Acc.
Unseen
Role Acc.
Exist
Mismatch
Acc.
Broken
Role Acc.
Random
Role Acc.
LLM Repurposed Llama 3.1 8B Instruct basic 42 84.11 16.50 15.00 85.54 86.33 81.89 78.00 43.00 100.00
LLM + LLM Repurposed Llama 3.1 8B Instruct basic 42 96.22 6.00 2.00 96.65 95.11 97.33 92.00 14.00 100.00
BERT + LLM Repurposed Modern BERT-base basic 42 90.56 14.25 5.60 91.74 91.89 89.22 81.00 53.00 100.00
LLM Repurposed Llama 3.1 8B Instruct office 42 84.56 22.50 11.00 86.65 87.89 80.11 70.00 49.00 100.00
LLM + LLM Repurposed Llama 3.1 8B Instruct office 42 88.11 20.25 4.00 89.86 89.22 87.00 73.00 17.00 100.00
BERT + LLM Repurposed Modern BERT-base office 42 87.89 21.75 4.40 89.77 88.78 87.00 71.00 33.00 100.00
LLM Synthetic Llama 3.1 8B Instruct basic 42 95.78 5.75 2.00 96.23 94.67 96.89 94.00 8.00 95.00
LLM + LLM Synthetic Llama 3.1 8B Instruct basic 42 98.11 2.25 2.00 98.30 98.11 98.11 97.00 7.00 100.00
BERT + LLM Synthetic Modern BERT-base basic 42 96.00 4.50 3.60 96.40 94.78 97.22 94.00 41.00 100.00
LLM Synthetic Llama 3.1 8B Instruct office 42 84.00 30.50 5.00 86.91 85.11 81.78 63.00 14.00 89.00
LLM + LLM Synthetic Llama 3.1 8B Instruct office 42 83.78 33.75 2.00 87.01 84.89 82.67 55.00 16.00 100.00
BERT + LLM Synthetic Modern BERT-base office 42 77.22 47.25 3.20 82.52 78.22 76.22 37.00 28.00 100.00
Average 88.86 18.77 4.98 90.63 89.58 87.95 75.42 26.92 98.67
Table 10: Access control metrics for Single Name Encoding
Architecture Dataset
Model
Org.
Structure
Seed Accuracy FPR FNR F1
Seen
Role Acc.
Unseen
Role Acc.
Exist
Mismatch
Acc.
Broken
Role Acc.
Random
Role Acc.
LLM Repurposed Llama 3.1 8B Instruct basic 42 90.44 11.25 15.00 91.43 90.44 90.44 78.00 43.00 100.00
LLM + LLM Repurposed Llama 3.1 8B Instruct basic 42 94.11 9.75 2.00 94.83 95.22 93.00 92.00 14.00 100.00
BERT + LLM Repurposed Modern BERT-base basic 42 93.44 10.50 5.60 94.24 94.00 92.89 81.00 53.00 100.00
LLM Repurposed Llama 3.1 8B Instruct office 42 85.56 18.75 11.00 87.25 87.78 84.44 70.00 49.00 100.00
LLM + LLM Repurposed Llama 3.1 8B Instruct office 42 88.33 18.75 4.00 89.95 90.56 86.11 73.00 17.00 100.00
BERT + LLM Repurposed Modern BERT-base office 42 88.89 17.50 4.40 90.38 89.44 88.33 71.00 33.00 100.00
LLM Synthetic Llama 3.1 8B Instruct basic 42 96.33 6.00 2.00 96.75 95.22 97.44 94.00 8.00 95.00
LLM + LLM Synthetic Llama 3.1 8B Instruct basic 42 98.56 2.25 2.00 98.71 98.56 97.44 97.00 7.00 100.00
BERT + LLM Synthetic Modern BERT-base basic 42 96.56 4.25 3.60 96.91 97.33 95.78 94.00 41.00 100.00
LLM Synthetic Llama 3.1 8B Instruct office 42 80.78 34.50 5.00 84.32 83.00 78.56 63.00 14.00 89.00
LLM + LLM Synthetic Llama 3.1 8B Instruct office 42 78.11 46.50 2.00 83.23 79.22 77.00 55.00 16.00 100.00
BERT + LLM Synthetic Modern BERT-base office 42 79.33 42.75 3.20 83.91 81.44 77.22 37.00 28.00 100.00
Average 89.20 18.56 4.98 90.99 90.19 88.22 75.42 26.92 98.67
Table 11: Access control metrics for Hierarchical Name Encoding
Architecture Dataset
Model
Org.
Structure
Seed Accuracy FPR FNR F1
Seen
Role Acc.
Unseen
Role Acc.
Exist
Mismatch
Acc.
Broken
Role Acc.
Random
Role Acc.
LLM Repurposed Llama 3.1 8B Instruct Basic 42 75.00 24.00 25.00 79.00 78.00 72.00 76.00 74.00 100.00
LLM + LLM Repurposed Llama 3.1 8B Instruct Basic 42 79.00 16.00 24.00 82.00 81.00 77.00 84.00 64.00 100.00
BERT + LLM Repurposed Modern BERT-base Basic 42 92.25 13.33 4.40 93.91 91.25 93.25 86.67 65.00 100.00
LLM Repurposed Llama 3.1 8B Instruct Office 42 80.00 27.00 15.00 84.00 84.00 77.00 73.00 49.00 99.00
LLM + LLM Repurposed Llama 3.1 8B Instruct Office 42 87.00 26.00 5.00 90.00 89.00 84.00 74.00 31.00 100.00
BERT + LLM Repurposed Modern BERT-base Office 42 86.75 27.33 4.80 89.98 89.00 84.50 72.67 50.00 100.00
LLM Synthetic Llama 3.1 8B Instruct Basic 42 89.00 19.00 7.00 91.00 89.00 89.00 81.00 62.00 95.00
LLM + LLM Synthetic Llama 3.1 8B Instruct Basic 42 97.00 3.00 2.00 98.00 98.00 97.00 97.00 43.00 100.00
BERT + LLM Synthetic Modern BERT-base Basic 42 89.75 17.33 6.00 91.98 88.50 91.00 82.67 71.00 100.00
LLM Synthetic Llama 3.1 8B Instruct Office 42 76.00 54.00 6.00 83.00 78.00 74.00 46.00 34.00 94.00
LLM + LLM Synthetic Llama 3.1 8B Instruct Office 42 81.00 48.00 2.00 87.00 83.00 79.00 52.00 20.00 100.00
BERT + LLM Synthetic Modern BERT-base Office 42 80.38 39.33 7.80 85.45 81.50 79.25 60.67 54.00 99.00
Average 84.43 26.19 9.08 87.94 85.85 83.08 73.81 51.42 98.92
Table 12: Access control metrics for Hierarchical Number Encoding
Struct. Arch. Model Acc. FPR FNR F1 Corr. Comp. Clar. Seen Unseen Repurposed Dataset (Dolly) Basic LLM Qwen2.5-3B 76.33 22.67 24.33 80.00 3.92 3.53 4.65 80.00 72.67 Basic LLM Llama-3.2-3B 76.33 28.00 21.67 80.33 4.02 3.64 4.65 78.00 74.00 Basic LLM gemma-4B 75.33 27.33 23.00 79.67 3.99 3.55 4.59 78.33 72.33 Basic LLM Qwen2.5-7B 76.33 24.33 24.00 80.00 4.08 3.67 4.71 78.67 73.33 Basic LLM Llama-3.1-8B 75.67 25.00 24.00 79.33 4.12 3.65 4.69 78.00 73.00 Basic LLM gemma-7B 73.00 34.00 22.33 78.33 3.85 3.55 4.45 76.00 70.33 Basic LLM-Cls Qwen2.5-3B 90.33 16.00 5.67 92.67 90.67 90.67 Basic LLM-Cls Llama-3.2-3B 89.00 18.00 6.67 91.67 90.67 88.00 Basic LLM-Cls gemma-4B 91.33 14.67 5.33 93.33 92.67 90.33 Basic LLM-Cls Qwen2.5-7B 85.67 22.67 9.33 89.00 86.67 85.00 Basic LLM-Cls Llama-3.1-8B 77.33 29.67 18.33 81.67 78.67 76.00 Basic LLM-Cls gemma-7B 78.33 37.33 12.67 83.33 78.67 77.67 Basic Cls Modern BERT-base 92.96 11.44 4.40 94.44 92.92 93.00 Basic Cls Modern BERT-large 92.58 12.11 4.60 94.15 92.75 92.42 Basic Cls Google BERT-base 86.82 29.33 1.97 90.36 88.00 86.43 Basic Cls Google BERT-large 75.77 56.61 4.69 82.95 75.77 77.28 Basic Cls RoBERTa-base 74.21 57.18 3.29 82.50 80.07 71.66 Basic Cls RoBERTa-large 85.83 16.45 9.99 89.54 85.50 86.49 Office LLM Qwen2.5-3B 76.67 25.33 22.33 80.33 3.93 3.47 4.63 80.67 72.67 Office LLM Llama-3.2-3B 83.00 25.33 11.67 86.67 3.99 3.59 4.66 86.00 80.00 Office LLM gemma-4B 79.33 25.67 17.67 83.33 4.08 3.72 4.59 81.67 77.33 Office LLM Qwen2.5-7B 80.00 25.67 16.33 84.00 4.19 3.74 4.73 84.00 76.00 Office LLM Llama-3.1-8B 80.33 26.67 15.00 84.33 4.17 3.70 4.68 83.67 77.33 Office LLM gemma-7B 80.00 24.67 17.67 83.67 3.77 3.41 4.40 83.67 75.67 Office LLM-Cls Qwen2.5-3B 86.67 27.67 4.67 89.67 88.33 84.33 Office LLM-Cls Llama-3.2-3B 88.67 22.00 5.33 91.00 89.67 87.33 Office LLM-Cls gemma-4B 86.33 27.00 5.33 89.67 88.33 84.33 Office LLM-Cls Qwen2.5-7B 87.00 26.33 5.00 90.33 89.33 85.00 Office LLM-Cls Llama-3.1-8B 86.33 28.33 4.67 90.00 88.67 84.00 Office LLM-Cls gemma-7B 87.67 24.67 4.67 90.33 89.33 86.00 Office Cls Modern BERT-base 86.38 25.22 6.67 89.52 87.67 85.08 Office Cls Modern BERT-large 87.38 25.67 4.80 90.41 88.75 86.00 Office Cls Google BERT-base 85.11 30.20 6.09 90.17 89.19 83.62 Office Cls Google BERT-large 86.96 29.65 6.34 91.12 89.29 85.03 Office Cls RoBERTa-base 83.15 27.18 9.83 85.68 85.35 84.16 Office Cls RoBERTa-large 63.75 99.72 0.81 77.13 62.85 62.41
Table 13: Role-aware performance on repurposed (Dolly) dataset. Green cells mark the best accuracy in each dataset block. Higher is better for all metrics except FPR/FNR (lower is better).
Struct. Arch. Model Acc. FPR FNR F1 Corr. Comp. Clar. Seen Unseen Synthetic Dataset Basic LLM Qwen2.5-3B 72.00 37.33 22.33 77.67 3.96 3.69 4.74 72.67 71.33 Basic LLM Llama-3.2-3B 92.00 12.67 5.33 93.33 3.86 3.60 4.68 91.33 92.33 Basic LLM gemma-4B 75.33 42.00 14.33 81.33 3.96 3.63 4.62 75.33 75.00 Basic LLM Qwen2.5-7B 77.33 35.00 15.33 82.00 4.04 3.78 4.78 79.00 75.00 Basic LLM Llama-3.1-8B 92.67 13.33 4.67 94.00 3.95 3.73 4.79 92.67 92.00 Basic LLM gemma-7B 78.33 34.33 14.67 83.00 3.93 3.66 4.62 79.67 76.00 Basic LLM-Cls Qwen2.5-3B 90.67 14.67 6.67 92.33 89.33 91.67 Basic LLM-Cls Llama-3.2-3B 96.67 5.33 2.33 97.33 96.33 97.00 Basic LLM-Cls gemma-4B 97.33 4.00 2.33 97.67 97.00 97.33 Basic LLM-Cls Qwen2.5-7B 96.33 6.33 2.00 97.00 96.33 96.00 Basic LLM-Cls Llama-3.1-8B 97.00 3.67 2.00 98.00 97.33 97.33 Basic LLM-Cls gemma-7B 91.67 19.33 2.00 93.67 91.67 91.67 Basic Cls Modern BERT-base 91.08 12.00 7.07 92.88 89.92 92.25 Basic Cls Modern BERT-large 84.50 25.33 9.60 87.92 84.25 84.75 Basic Cls Google BERT-base 87.48 25.74 3.05 90.96 86.80 90.83 Basic Cls Google BERT-large 90.73 15.81 5.90 92.94 90.95 91.23 Basic Cls RoBERTa-base 80.67 48.27 3.67 85.95 80.46 80.61 Basic Cls RoBERTa-large 61.45 74.25 12.94 74.83 62.30 66.95 Office LLM Qwen2.5-3B 77.67 47.67 7.00 83.67 3.76 3.60 4.71 80.00 75.67 Office LLM Llama-3.2-3B 78.67 47.33 5.67 84.67 3.85 3.71 4.73 80.33 77.00 Office LLM gemma-4B 73.67 58.00 7.33 81.67 3.84 3.69 4.69 76.33 71.00 Office LLM Qwen2.5-7B 79.00 45.33 6.33 84.67 3.89 3.77 4.77 81.00 77.00 Office LLM Llama-3.1-8B 78.00 49.00 6.00 84.00 3.94 3.77 4.81 80.00 76.00 Office LLM gemma-7B 76.00 53.33 6.33 83.00 3.81 3.59 4.58 80.00 71.67 Office LLM-Cls Qwen2.5-3B 79.67 51.33 2.00 85.67 81.00 78.33 Office LLM-Cls Llama-3.2-3B 80.00 50.00 2.00 85.67 81.00 79.00 Office LLM-Cls gemma-4B 79.67 51.00 2.00 85.33 81.67 77.67 Office LLM-Cls Qwen2.5-7B 81.33 45.33 2.33 86.67 82.33 80.33 Office LLM-Cls Llama-3.1-8B 81.67 46.67 2.00 87.00 84.00 79.00 Office LLM-Cls gemma-7B 80.00 49.33 2.00 86.00 81.33 79.00 Office Cls Modern BERT-base 80.17 43.89 5.40 85.63 82.08 78.25 Office Cls Modern BERT-large 77.13 53.33 4.60 83.91 78.17 76.08 Office Cls Google BERT-base 75.32 62.32 3.97 83.51 78.18 73.29 Office Cls Google BERT-large 78.17 55.27 4.67 85.57 77.10 77.62 Office Cls RoBERTa-base 73.79 63.95 3.63 82.71 76.38 72.93 Office Cls RoBERTa-large 69.13 79.94 0.73 80.69 70.98 69.10
Table 14: Role-aware performance on synthetic datasets. Green cells mark the best accuracy in each dataset block. Higher is better for all metrics except FPR/FNR (lower is better).